Determine the approximate freezing point of 0.1 F potassium sulfate solution
1
Expert's answer
2020-01-28T06:14:48-0500

"\Delta T_f=" "i*" "m*K_f" ...............(1)

where ΔTf is the freezing point depression,i is the vanhoff's factor, m is the molality of the solute and Kf is the freezing point depression constant

here value of Kf for water is 1.86 , m is 0.1

now calculating value of i

As potassium sulfate solution will undergo 100 % dissociation

"K_2SO_4\iff2K^++SO_4^{2-}"

so total 3 ions are formed on 100 % dissociation so value of i =3


Substituting these values in the equation (1)

we get "\Delta T_f=3*0.1*1.86=0.558 \ ^0C"

freezing point ="T_f -\Delta T_f" ="(0-0.558)^0C=-0.558^0C"